40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 166 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  9163   Thu Sep 26 01:49:28 2013 MasayukiSummaryGreen LockingFPMI noise caused by ARM locking

 

 Measurement with FPMI


i)By locking the FPMI with AS55Q and arms using POX,POY we measured  the OLTF on AS55Q, the response from BS actuation to error signal on AS55Q  for H_mich. The fitted,  measured OLTF and the residual function is in attachment1. I fitted two parameters and they are time-delay and the gain. The time delay is -275 usec. The time delay in three different control are almost same. The response from BS to AS55Q is in attachment 2.


With these two measuremets, I calclated the H_mich in FPMI. This H_mich should be different from simple MI because the cavity  refrectivity is different from the front mirror. Acrually it changed and the value was
Hmich = 4.4026e7

ii) I excited the ETMX and ETMY and measure the response from actuation to the error signal of MICH on AS55Q. The response is in attachment 3 and 4. from these result I calculated the H_L-l by using the formula as I mentioned. The value was
H_Lx-l = 175.7650 (XARM)
H_Ly-l = 169.8451 (YARM)

iii) I measured the error signal of MICH and XARM and YARM and with measured H_L-l, I estimated the FPMI noise caused by ARM locking. You can see in the higher frequency region than 10 Hz is dominated by noise caused by ARM control in-loop noises. 150 Hz and 220Hz are the UGF of each arms, so the two peaks are caused by arm control. You can see the small difference between FPMI noise and  noise from arms. There are two possibilities, one is that these measurement is not same time measurement so they should have small difference. and  other possibility is the error of the caliculation. But I think it doesn't look so bad estimation.

 

Next step

We will do same measurement with lock the arms the ALS system on tomorrow. Then we will check the PDH servo or other noise source and investigate the ALS system

Attachment 1: MICH_OLTF.pdf
MICH_OLTF.pdf
Attachment 2: BS-RS55Q.png
BS-RS55Q.png
Attachment 3: ETMX-RS55.png
ETMX-RS55.png
Attachment 4: ETMY-RS55.png
ETMY-RS55.png
Attachment 5: plot.pdf
plot.pdf
  9162   Wed Sep 25 23:59:29 2013 MasayukiSummaryGreen LockingFPMI noise caused by ARM locking

 

Measurement with ARMs

i) By locking the MICH with AS55Q signal I measured the actuator response of ITMX ITMY BS for calibration of each actuator. This measurement was done at the same time with elog#9158. The actuator response was
 
BS : 2.2347e-8 / f^2 [m/count]

ITMX: 5.0843e-9 /f^2 [m/count]

ITMY: 4.9677e-9 / f^2 [m/count]

 
 
ii)By locking the Arms for IR with POX,POY. I measured the OLTF and the response from ITM actuation to POX and POY signal. Attachment 1,2 are the plots of fitted OLTF , the measured OLTF, and residual function (model - measure)/model and the attachment 3,4 are the response of each arm. I fitted the three parameters. Those are the gain, time-delay and cavitypole. Each fitted parameter is
 
XARM ;
timedelay:-282.09 usec, cavity pole : 2872.0 Hz
YARM ;
timedelay:-283.84 usec, cavity pole : 2939.9 Hz
 
The cavity pole seems higher than privious measurement (In 2009). Actually the residual function are increase at the higher frequency region than 1kHz, so I guess the fitting is not so good.One possibility is that in the region where cavity pole effect increase we has not much data.
With fitted OLTF and actuator response I calibrated the H_xarm and H_yarm.
 
Hxarm : 2.9796 e11 [count / m]
Hyarm : 6.1394 e11 [count / m]
 
iii) After that I measured the response from ETM actuation to POX and POY signal to calibrate the ETM actuator. The response of each actuator is
 
ETMX:1.2040e-8 / f^2 [m/count]

ETMY:1.4262e-8 / f^2 [m/count]
 
iv) I calibrated the error signal with OLTF and Hxarm,Hyarm. The result is in Attachment 5

 In high frequency region there is the difference between xarm and yarm. These difference are already there in error signal. I'm not sure where these noise comes from. We will make measurement with Green PDH from tomorrow, so  we can also check with those measurement.

In other region the two noises are very close and also very similar to the plot of the seismic motion in the control room (attached on the front of TV screen).

Attachment 1: XARM_OLTF.pdf
XARM_OLTF.pdf
Attachment 2: YARM_OLTF.pdf
YARM_OLTF.pdf
Attachment 3: XARM_ITMXresponse.png
XARM_ITMXresponse.png
Attachment 4: YARM_ITMYresponse.png
YARM_ITMYresponse.png
Attachment 5: free_running.pdf
free_running.pdf
  9161   Wed Sep 25 23:15:11 2013 MasayukiSummaryGreen LockingFPMI noise caused by ARM locking
I measured some error signal, OLTFs and responses for FPMI noise estimation. Especially we are interested in the noise from in-loop noise of ALS Green PDH control. The strategy and
 
1) Purpose
 Estimation of the FPMI phase shift noise caused by in-loop noise of Green PDH control. 
 
 
2) What we should figure out
 For that estimation we have to figure out the transfer function from the cavity length change to the phase shift which is measured by MICH.
 
 
3) Strategy
 I attached the block diagram of  our interferometer. Our goal is to find the transfer function H_L-l and to calibrate the out of loop noise of interferometer with that TF and error signal of the PDH control.
 H,A and F mean the sensitivity, actuator response and servo filter for each control loop. L_xarm is the disturbance of the cavity length and l- is the differencial motion of the interferometer
fpmibd.pdf
We can get this H_L-l from measurement of the response from calibrated ETM actuation to the MICH error signal. You can get the formula for calculating H_L-l with simple calculation and that is
 
             1 + G_mich       1 + G_xarm      V_mi  
H_L-l = ---------------  -----------------  ------------
              H_mich             A_etmx         V_excetm
 
 
where the each G is OLTF and V_mi/Vexcetm is the response from the ETM actuation to the MICH error signal.
And then  the FPMI noise in the unit of meter / rHz is
 
                           H_L-l
N_fpmi = l_dis + ------------ Vx
                          H_mich
               
This second term is what we are interested in.
 
To estimate these noises
i) We can calibrate the actuators of  ITMX, ITMY and BS with using the MICH as sensor. So we can calibrate the arm error signals by  excitation of arm length using ITMs actuator.
ii) If we know the TFs of arms, we can calibrate the ETMX and ETMY actuators.
iii) We should know the response from ETMX or ETMY actuating to error signal of mich.
iv) Also we should calibrate the error signal of MICH in FPMI locking(H_mich). We can do that by exciting the BS.
 

Then we can estimate the noises.

 
In next entry, I will write about measurement.

 

  9160   Wed Sep 25 19:34:51 2013 ranaUpdateSUSProblems with ETMY Optical Lever

I went down to investigate the issue with the extra noise that I found in the ETMY optical lever yesterday. There were several problems with the optical layout down there - I'm not sure if I remember them all now.

  1. Beam reflected from OL QPD not dumped.
  2. OL QPD set normal to the steering mirror so that the back reflection goes into the vacuum chamber.
  3. HeNe laser mount only dogged with 2 dogs. Needs 3. Looks like some said "Aw, that's not goin' nowhere. Let's just leave that there pard!"
  4. First lens downstream of the laser had 2 screws and washers, but neither was even finger tight! They were loose by more than 1 full turn.
  5. Second lens was clipping. Beam was so far off center that this lens was being used to steer the beam by a few inches on the QPD.
  6. Extra reflections from ingoing beam (I don't know which surfaces) randomly landing on green & red optics.
  7. Lenses for the HeNe mode matching are coated for 1064 nm. HeNe is 633 nm, so these lenses must be replaced to reduce the reflections.

The main noise issue, however, appears to be not a layout issue at all. Instead its that the laser intensity noise has gone through the roof. See attached spectra of the quadrants (this is the way to diagnose this issue).

I'll ask Steve to either heal this laser or swap it out tomorrow. After that's resolved we'll need another round of layout fixing. I've done a couple of hours today, but if we want a less useless and noisy servo we'll have to do better.

NOTE: by looking at the OL quadrants, I've found a noisy laser, but this still doesn't explain the excess noise in the ETMX. That was the one that has a noisier error signal, not ETMY. By the coherence in the DTT, you can see that the ETMY OL is correctly subtracting and normalizing out the intensity noise of the laser. Seems like the ETMX electronics might be the culprit down there.

Attachment 1: ETMY-BadHeNe.pdf
ETMY-BadHeNe.pdf
  9159   Wed Sep 25 17:07:08 2013 ranaFrogsTreasureFree Green Mango Juice in fridge

aam_pana_recipe.jpg

its an acquired taste, but its a must since we're sending an interferometer to India

  9158   Wed Sep 25 08:11:01 2013 MasayukiSummaryLSCLSC calibration screen

 

  The real time calibration system is not correct in high frequency.

The attachment are the plot of two free running noise. Blue curve is the plot of noise calibrated with OLTF. Green one is the just fft analysed signal of the real time calibration system output. You can see the ripple in high frequency region in green curve. That is because the anti-aliasing filter and digital anti-aliasing filter. I assume the sensitivity of MI as constant but Rana mentioned that we should take these filters into account.

modeled OLTF and sensitivity H
 I put the AA filter and DAA filter effects into matlab calibration script. The attachment 2 is the modeled sensitivity of the MICH. You can find each filter properties in  elog#8555 (analog AA filter) and in elog#3961. I estimate the H gain by measuring the fringe. The attachment 3 is the plot of fringe and I averaged with green points. The actual number is 3.48e7 count/m

attachment 2: the sensitivitiy of MICH

Screenshot.png

 attachment 3: fringe of the MICH

fringe.pdf

I modeled OLTF with this H and the fitted into the measurement data. That is in attachment 4. In this OLTF I also included the DAI filter and AI filter, and ' sample and hold circuit' of DAC TF . These are  mentioned in two references. Additionally I added the time delay 309.6 us.  Yuta mentioned that in C1SUS has 125us time delay. In MICH control we have also C1LSC , so I think this time delay is reasonable. I compensated the error signal with these OLTF and MICH sensitivity.

attachment 4: OLTF of the MICH control

Screenshot-1.png

You can see that the ripple is gone in blue curve and after 5 kHz the curve is flat.

Next step

I'm trying to put the inverted AA filter and DAA filter in C1CAL_INCV servo. But the ploblem is the difference of sampling frequency, so I couldn't fix yet. One possibility is putting approximated filter. I hope I will find some good way to design these filters.

 

 

Other thing

I esitimated the FPMI noise propagated from the residual noise of IR PDH control of both ARMS. I will summarize and write these staff in this afternoon.

Attachment 1: plot.pdf
plot.pdf
  9157   Tue Sep 24 22:19:57 2013 ManasaUpdateGeneralFPMI locked

 [Masayuki, Manasa]

We locked FPMI and measured the FPMI noise (power spectrum of error signal - MICH_IN1) which will be calibrated.

The arms were locked using POX11 and POY11. The sign of MICH gain was changed to lock FPMI (from -30 to +30). 

  9156   Tue Sep 24 20:43:45 2013 masayukiSummarySUSoptical levers centering

I centered optical levers of ITMX,BS,ETMY. I also change the position of optical levers of ITMX, ETMY, ITMY, BS on Friday night(9/21), of ITMX, ETMY, BS on Monday night. Both are around 6:00 ~ 7:00.But centering on Monday was totally wrong, because I centered with not good IFO alignment.

The attachment is the 5 days trend of the opt lev of ITMX. First gap is alignment on Friday and Second gap is the alignment on Monday. Now I centered after  locking the FPMI.

The attachment 2 is the last 6 hours data.  The gap on 9/25 00:00 and 1:30(UTC)  is because the alignment of the cavity and the last gap is because of  centering of the optical lever.

Attachment 1: Screenshot-Untitled_Window.png
Screenshot-Untitled_Window.png
Attachment 2: Screenshot-Untitled_Window-1.png
Screenshot-Untitled_Window-1.png
  9155   Tue Sep 24 10:55:45 2013 ranaUpdatePSLPMC re-aligned

After relocking the PMC at a good voltage, Steve and I re-aligned the beam into the PMC by walking the last two steering mirrors. After maximizing the power, we also aligned the reflected beam by maximizing the PMC_REFL_DC with the unlocked beam.

Transmission is back to 0.84 V. We need Valera mode matching maintenance to get higher I guess. Maybe we can get a little toaster to keep the PMC PZT more in the middle of its range?

Attachment 1: psl-trend.png
psl-trend.png
  9154   Sun Sep 22 23:04:52 2013 ranaUpdatePEMGuralp needs recentering

 After seeing all of these spikes in the BLRMS at high frequency for awhile, I power cycled the Guralp interface box (@ 10:21 PM) to see if it would randomly recenter in a different place and stop glitching.

It did - needs to be better centered (using the paddle). Plot shows how the Z channel gets better after power cycle.

Attachment 1: seis.pdf
seis.pdf
  9153   Sun Sep 22 22:54:28 2013 ranaUpdateIOOmode cleaner not locking

Having trouble again, starting around 1 hour ago. No one in the VEA. Adjusted the offset -seems to be OK again.

  9152   Sun Sep 22 22:05:10 2013 ranaUpdateSUSoplev XY-plots reflect new calibration

The ETMX oplev signal looks kind of dead compared to the ETMY. It has no features in the spectra and the SUM is pretty low.

I noticed that the cal fields are still set to 1. To get it close to something reasonable, I calibrated it vs. the SUSPIT and SUSYAW values by giving it a step in angle and using 'tdsavg' plus some arithmetic.

OLPIT =  45 urads/ count

OLYAW = 85 urads / count

These are very rough. I don't even know what the accuracy is on the OSEM based calibration, so this ought to be redone in the way that Jenne and Gabriele did before.

The attached image shows the situation after "calibration" of ETMX. This OL system needs some noise investigation.

Attachment 1: noise.png
noise.png
  9151   Sun Sep 22 21:28:53 2013 ranaUpdateSUSset OL T RAMP values (they are not visible on the OL screens)

controls@rosalba:/opt/rtcds/caltech/c1/scripts/SUS 0$ ./setOLtramps
Old : C1:SUS-ETMX_OLPIT_TRAMP        0
New : C1:SUS-ETMX_OLPIT_TRAMP        2
Old : C1:SUS-ETMX_OLYAW_TRAMP        0
New : C1:SUS-ETMX_OLYAW_TRAMP        2
Old : C1:SUS-ETMY_OLPIT_TRAMP        2
New : C1:SUS-ETMY_OLPIT_TRAMP        2
Old : C1:SUS-ETMY_OLYAW_TRAMP        2
New : C1:SUS-ETMY_OLYAW_TRAMP        2
Old : C1:SUS-ITMX_OLPIT_TRAMP        0
New : C1:SUS-ITMX_OLPIT_TRAMP        2
Old : C1:SUS-ITMX_OLYAW_TRAMP        0
New : C1:SUS-ITMX_OLYAW_TRAMP        2
Old : C1:SUS-ITMY_OLPIT_TRAMP        0
New : C1:SUS-ITMY_OLPIT_TRAMP        2
Old : C1:SUS-ITMY_OLYAW_TRAMP        0
New : C1:SUS-ITMY_OLYAW_TRAMP        2
Old : C1:SUS-BS_OLPIT_TRAMP          0
New : C1:SUS-BS_OLPIT_TRAMP          2
Old : C1:SUS-BS_OLYAW_TRAMP          0
New : C1:SUS-BS_OLYAW_TRAMP          2
Old : C1:SUS-PRM_OLPIT_TRAMP         0
New : C1:SUS-PRM_OLPIT_TRAMP         2
Old : C1:SUS-PRM_OLYAW_TRAMP         0
New : C1:SUS-PRM_OLYAW_TRAMP         2
Old : C1:SUS-SRM_OLPIT_TRAMP         0
New : C1:SUS-SRM_OLPIT_TRAMP         2
Old : C1:SUS-SRM_OLYAW_TRAMP         0
New : C1:SUS-SRM_OLYAW_TRAMP         2
 
Done setting TRAMPs

  9150   Sun Sep 22 21:03:15 2013 ranaConfigurationSUSTuned bounce and roll mode of ETMY suspension

 Today I noticed that there was a lot of noise at the Bounce and Roll eigenfrequencies for ETMY. I found that the bandstop filter were set at completely the wrong frequencies, so I've remade them.

The filters were last tuned by Leo in May of 2011. Even so, he left the frequencies at the frequencies of the old MOS suspensions which had f_bounce ~ 12 Hz.

 The FOTON plot shows the OLD ones versus the NEW ones. The DTT spectra shows the oplev error signals in the usual state. I have also copied these over to the SUSPOS,PIT,YAW, and SIDE filter banks and turned them all ON.

I also turned OFF and deleted the 3 Hz RG filter that was there. There's no such peak in the error signal and even if one wanted to compensate for the stack mode, it should be a low Q filter, not this monster.

Attachment 1: etmy-error.png
etmy-error.png
Attachment 2: notches.pdf
notches.pdf
  9149   Fri Sep 20 22:49:10 2013 MasayukiUpdateLSCLSC calibration screen

Quote:

I update the LSC calibration screen. This screen is for real time calibration of each DOF with using error signal and control signal. The formula of the calibration is

x_dis = V_err/H + A V_fb

,where x_dis is the disturbance without surpression, V_err and V_fb are error signal and control signal, H is the transfer function from the displacement to output and A is the efficiency of the actuator.

I will put the filter of 1/H into the CINV filter bank and actuator efficiency into the A filter bank.

 

 I fixed the filter of the MICH real-time calibration. You can find C1CAL screen from the LSC menu 'calibration' of sitemap.

*Filter explanation

    C1CAL_MICH_CINV : the servo to convert  the error signal to displacement.

Sen_MICH :

the inverse of the transfer function from the distance to the error signal, which has the unit of count/m. In the formula this filter is represented by 1/H.

I assume this H is independent of frequency and time, and I calculated by the amplitude of the fringe of error signal. But it may change every day by drift of laser intensity and so on.  So we should follow the actual H somehow.  The temporary value of H is 3.76*10^7 count/m .

    C1CAL_MICH_A : the servo to convert the feedback signal to displacement. In formula This transfer function is represented by A

 SUS_BS;

the transfer function of the suspension of the BS. This is modeled from the measurement in elog#9127. The resonant frequency is 1.029 Hz and Q is 12.25.

 Res_A :

the response of the actuator on BS_SUS, which has the unit of m/count. The value is 1.99*10^-8 m/count. This value is measured in the measurement in elog#9121.

     C1CAL_MICH_W : the servo to handle the calibrated signal.

  m->um ;

the filter to convert the unit of signal from m to um. When this filter is on, the output is written in unit of um.

*Measurement

 I measured the power spectrum of the calibrated free running noise. The measured port was C!CAL_MICH_W_OUT. The result is in attachment 1. Also in this figure there are the plots of the Verr/H and Vfb*A.

 In low frequency region, where control loop suppresses the disturbance, you can see that the displacement is equal to the displacement of actuation (I'm not sure what happens at the point of 0.03Hz), and in high frequency region, where control loop doesn't work, the displacement is equal to the value of the Verr divided by MICH sensitivity. Also this result is similar to the my calibration result.elog#9131

Attachment 1: plot.pdf
plot.pdf
  9148   Fri Sep 20 20:27:18 2013 ranaUpdateIOOmode cleaner not locking

 I used our procedure from this entry to set the IMC board offset as well as the FSS board offset.

I found this afternoon that the MC was having trouble locking: the PC path was railing as soon as the boost was engaged. Could be that there's some misalignment on the PSL which has led to some RAM having to be canceled by this new offset. Let's see if its stable for awhile.

  9147   Fri Sep 20 20:14:52 2013 ranaSummaryGeneral/home/cds nearly full

Quote:

One of the reasons that our disk is getting full is due to the scripts_archive directory. A backup script runs on op340m and makes a tar.bz2 file of the scripts directory and puts it in scripts_archive every morning at 6 AM.

On Oct 7, 2011, Koji fixed this script to point at our new scripts directory instead of the old /cvs/cds/caltech/scripts directory. Since then, however, no one has fixed the exclude file to NOT back up the junk that's in that directory. Its a 1.6 GB directory so its full of it.

I've deleted a bunch of junk from the scripts directory: this directory is for scripts, not for your personal home movies or junk data files. Put those in your USER directory. Put temporary data files in /tmp/. I've also added a few more patterns to the exclude file so that less .mpg, .png, .pdf, .dat, etc get stored every day. The new daily .tar.bz2 file wil be ~25 MB instead of 770 MB.

(also fixed the backup script to use 'env' to setup the perl environment and removed the hard-coded path to tar)

 OUr disk was getting full again. Turned out my "fix" to 25 MB was only a fix to 250 MB. Since we were getting disk full warnings on our Ubuntu workstations, I deleted some COMSOL.dmg files from users/zach/ and then started deleting every other tarball from the scripts_archive directory. ~221 GB are now free. Still need to fix the exclude file for scripts better.

  9146   Fri Sep 20 16:37:30 2013 SteveUpdateVACvertex crane folding arm is back on

Quote:

 The folding beam removed as shown. Two man supporting it while I hammering it out. Pin was dry and it gulled into supporting hinges.

The rotating hinge will be machined and bushing will be added with Zerk fitting or similar. This will allow lubrication in the future.

 see elog #9111

Atm1,  The folding arm is back on with 0.1" misalignment at no load in the trolly's way. The other side of the I beam is 0.02" higher than the main beam.. New bushing and pin were greased up  with Krytox before installation.

The axial Zerk 1/8" pipe in the pin upper end can not take any fitting. There is no room. It is taped off.

This gap comes down to ~ 1/16" at fully extended arm with 225 lbs load at the end of it.

The present plan is to grind down the  the misalignment of 0.1"  for a slow-loaded trolly.

Steve Baker of Konacranes will be back to grind down this ridge and load test at 500 lbs on Tuesday,  OCT  1, 2013

 

 

 

Attachment 1: bushingBottom.jpg
bushingBottom.jpg
Attachment 2: foldingI.jpg
foldingI.jpg
Attachment 3: noLoadGap.jpg
noLoadGap.jpg
Attachment 4: trollyway.jpg
trollyway.jpg
Attachment 5: trollySee.jpg
trollySee.jpg
  9145   Fri Sep 20 09:49:06 2013 SteveUpdateSUSOptical Lever Trend for 180 days: bad ETMY & PRM

I'm working on it.

Attachment 1: ETMYoplevPRM.png
ETMYoplevPRM.png
  9144   Fri Sep 20 08:15:30 2013 SteveUpdateComputer Scripts / ProgramsOttavia, Rossa and Pianosa

Ottavia, Rossa  and Pianosa are running out of storage  space.

Attachment 1: Ottavia.png
Ottavia.png
Attachment 2: Rossa.png
Rossa.png
Attachment 3: Pianosa.png
Pianosa.png
  9143   Thu Sep 19 21:42:18 2013 ranaUpdateSUSOptical Lever Trend for 90 days: ETMX and PRM are the bad ones
Attachment 1: OLtrend_2013.png
OLtrend_2013.png
  9142   Thu Sep 19 21:15:44 2013 ranaUpdateComputer Scripts / ProgramsPMC locker

I used a script (~PSL/PMC/testAutoLocker.sh) to unlock the PMC and run autlocker ~100 times to see how robust the new autlocker is.

It failed to grab it 2 out of 137 times. During those times it just went on trying to ramp the PZT even after it had gone to a rail. Once someone resurrects Rob's 'trianglewave'  script we should be OK. Even so, I think this is good enough. Please try this out via the yellow button next time the PMC needs to be locked.

It usually takes 10-30 seconds to lock, depending upon where the fringe is compared to the upper voltage rail. Good enough.

Attachment 1: Untitled.pdf
Untitled.pdf
  9141   Thu Sep 19 18:48:24 2013 ranaUpdateComputer Scripts / ProgramsPMC locker

 In May of 2013 Den wrote a PMC Autolocker because he ignored / didn't want to read anyone else's code. Later that year Yuta also wrote another one from scratch for the same reasons.

I tried to use both today, but neither one runs. Yuta's one doesn't run because he was using a bunch of private yuta library stuff in the yuta directory. That kind of programming style is pretty useless for us since it never works after some time.

So I re-activated and tested the PMCAutolock bash script (it is actually a symbolic link called "PMCAutolock" which points to AutoLock.sh). These scripts are all basically the same:

They turn off the loop (or turn down the gain) and then scan the PZT, look for a resonance, and then activate the loop.

One problem with the logic has been that turning off the loop makes the gain so low that the peak flashes by too fast. But leaving the loop ON and just sweeping with the gain turned down to -10 dB is also not good. That only reduces the UGF from 1 kHz to ~100 Hz. What we want is more like a 10 Hz UGF while scanning the length. SO, I edited the script to turn down the modulation depth on the EOM by that factor. After acquiring lock, it returns all settings to the nominal levels as defined on the PSL_SETTINGS screen.

I've tested it a few times and it seems to work OK. You can run it from the yellow shabang button on the PMC screen.

I also changed the .bashrc aliases for the MEDM command so that if you type medm_good at the command line you get MEDM screens with scalable fonts. So you can stretch the screens.

Attachment 1: pmc.png
pmc.png
  9140   Thu Sep 19 10:24:58 2013 SteveUpdatePEMPRM damping restored after earth quake

Local m3.8 eq shakes PRM lose.

Attachment 1: eq3.8mLaVerne.png
eq3.8mLaVerne.png
  9139   Wed Sep 18 14:57:12 2013 SteveUpdateVACvertex crane folding arm removed

 The folding beam removed as shown. Two man supporting it while I hammering it out. Pin was dry and it gulled into supporting hinges.

The rotating hinge will be machined and bushing will be added with Zerk fitting or similar. This will allow lubrication in the future.

 see elog #9111

Attachment 1: manPower.jpg
manPower.jpg
Attachment 2: rotatingHinge.jpg
rotatingHinge.jpg
Attachment 3: supportingHinge.jpg
supportingHinge.jpg
Attachment 4: gullingPin.jpg
gullingPin.jpg
  9138   Wed Sep 18 11:52:53 2013 JamieUpdateCDSDataviewer cannot connect to fb

Quote:

Masayuki pointed out that dataviewer wasn't connecting to the fb this morning.

When I started dataviewer from the terminal I obtained the following error:

controls@pianosa:~ 0$ dataviewer
Can't find hostname `fb:8088'
Can't find hostname `fb:8088'; gethostbyname(); error=1
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Error in obtaining chan info.
Can't find hostname `fb:8088'
Can't find hostname `fb:8088'; gethostbyname(); error=1

I checked the CDS FE status screen and it looks normal. I could ping the fb and ssh to it as well.

I restarted fb to see if it made any difference. telnet fb 8088

It hasn't helped. Anything else that can be done??

I've fixed the problem.  This was due to a change I made in the NDSSERVER environment variable so that it would work with cdsutils.  I didn't realize there was an incompatibility with how dataviewer parses NDSSERVER.  Joe and I will have to figure it out.

In the mean time I've changed things back so that that dataviewer should now work as expected.  You might have to log out and back in for it to work (or at least open a new terminal).

  9137   Wed Sep 18 11:29:43 2013 manasaUpdateCDSDataviewer cannot connect to fb

Masayuki pointed out that dataviewer wasn't connecting to the fb this morning.

When I started dataviewer from the terminal I obtained the following error:

controls@pianosa:~ 0$ dataviewer
Can't find hostname `fb:8088'
Can't find hostname `fb:8088'; gethostbyname(); error=1
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Error in obtaining chan info.
Can't find hostname `fb:8088'
Can't find hostname `fb:8088'; gethostbyname(); error=1

I checked the CDS FE status screen and it looks normal. I could ping the fb and ssh to it as well.

I restarted fb to see if it made any difference. telnet fb 8088

It hasn't helped. Anything else that can be done??

CDS_FE.png

  9136   Tue Sep 17 18:44:29 2013 MasayukiUpdateLSCLSC calibration screen

I update the LSC calibration screen. This screen is for real time calibration of each DOF with using error signal and control signal. The formula of the calibration is

x_dis = V_err/H + A V_fb

,where x_dis is the disturbance without surpression, V_err and V_fb are error signal and control signal, H is the transfer function from the displacement to output and A is the efficiency of the actuator.

I will put the filter of 1/H into the CINV filter bank and actuator efficiency into the A filter bank.

 

Attachment 1: calibration_screen.png
calibration_screen.png
  9135   Tue Sep 17 17:55:42 2013 Jamie.ConfigurationComputer Scripts / Programspyepics configured

Quote:
 controls@rosalba:~ 0$ cdsutils
Traceback (most recent call last):
  File "/ligo/apps/cdsutils/lib/cdsutils/__main__.py", line 7, in <module>
    from cdsutils import CMDS
  File "/ligo/apps/cdsutils/lib/cdsutils/__init__.py", line 4, in <module>
    from servo import servo
  File "/ligo/apps/cdsutils/lib/cdsutils/servo.py", line 1, in <module>
    from epics import PV
ImportError: No module named epics
controls@rosalba:~ 1$
Mon Sep 16 19:40:32 2013 

 I properly installed the python-pyepics package on all the workstations, so this should be working now.

And for posterity, the pyepics source is at:

  pianos:/home/controls/src/pyepics

From this debian packages were built:

  controls@pianosa:~/src/pyepics 0$ debuild -uc -us

The .deb was then moved into the /ligo/apps/deb nfs:

  controls@pianosa:~/src 0$ cp python-pyepics_*_all.deb /ligo/apps/debs/pyepics/

It was then installed on the various workstations:

  controls@rosalba:~ 0$ sudo dpkg -i /ligo/apps/debs/pyepics/python-pyepics*.deb

This will probably need to be repeated any time we upgrade the EPICS install.

  9134   Tue Sep 17 00:50:42 2013 MasayukiUpdateLSCMICH calbration

I found the bug in my calibration code, and I fixed it.

And I put the white Gaussian noise on the BS actuator, and calibrated to the differential length with my new code. We already know the efficiency of the actuator(elog#8242), so I could estimate how much I put the disturbance and compare the two values. The result is in attachment 1.  x_exc means the value of the disturbance. 

You can see the PSD of the differential motion decrease factor of 3 by decreasing the disturbance by factor of 3 (except for the region from 1 Hz to 5 Hz), and the value at lower frequency than resonant frequency of the suspension is comparable to the value estimated with the actuator efficiency. Also there is no dip when I put the larger disturbance than free running noise.

Between 1 Hz and 5 Hz there seems to be a resonance of something (seismic stack?). And also on resonance of the suspension there seems to be some other noise source. One possibility is the active damping of each suspension.

Actually still there seems to be a dip between 0.1 Hz and 1 Hz. But if you consider about those effect, I think this result doesn't seems to be so strange. But according to the documentation of LIGO document-T000058, which I found the seismic motion in 40 m Lab is written in, the seismic motion at 0.1 Hz is 10^-7. I'm not sure about this factor of 10 difference. One possibility is the geophone doesn't have good sensitivity at low frequency. I'm still not sure this result is really collect.

 


 

Attachment 1: noise.PNG
noise.PNG
  9133   Mon Sep 16 19:41:01 2013 ranaConfigurationComputer Scripts / Programscdsutils checked out into /opt/rtcds
 controls@rosalba:~ 0$ cdsutils
Traceback (most recent call last):
  File "/ligo/apps/cdsutils/lib/cdsutils/__main__.py", line 7, in <module>
    from cdsutils import CMDS
  File "/ligo/apps/cdsutils/lib/cdsutils/__init__.py", line 4, in <module>
    from servo import servo
  File "/ligo/apps/cdsutils/lib/cdsutils/servo.py", line 1, in <module>
    from epics import PV
ImportError: No module named epics
controls@rosalba:~ 1$
Mon Sep 16 19:40:32 2013 
  9132   Mon Sep 16 15:29:50 2013 JamieConfigurationComputer Scripts / Programscdsutils checked out into /opt/rtcds

We now have a proper install of cdsutils:

 controls@rossa:~ 0$ cdsutils
 usage: cdsutils <cmd> <args>

 Advanced LIGO Control Room Utilites

 Available commands:

   read         read EPICS channel value
   write        write EPICS channel value
   switch       switch buttons in standard LIGO filter module
   avg          average NDS channels for some amount of time
   servo        simple integrator (pole at zero)

 Add '-h' after individual commands for command help.
 controls@rossa:~ 0$ 

It is installed in /ligo/apps/cdsutils, and should be in the path on all workstations.

The "development" source working directory is currently checked out at /opt/rtcds/cdsutils/trunk.

 

  9131   Mon Sep 16 14:11:47 2013 ranaUpdateLSCMICH calbration

  There doesn't seem to be any coherence among the different directions of ground motion (as expected from seismic theory), so I am suspicious of such a low MICH noise.

Attachment 1: Screen_Shot_2013-09-16_at_2.10.31_PM.png
Screen_Shot_2013-09-16_at_2.10.31_PM.png
Attachment 2: Screen_Shot_2013-09-16_at_2.18.47_PM.png
Screen_Shot_2013-09-16_at_2.18.47_PM.png
  9130   Mon Sep 16 13:11:15 2013 EvanUpdateComputer Scripts / ProgramsComsol 4.3b upgrade

Comsol 4.3b is now installed under /cvs/cds/caltech/apps/linux64/COMSOL43b. I've left the existing Comsol 4.2 installation alone; according to the Comsol installation guide [PDF], it is unaffected by the new install. On megatron I've made a symlink so that you can call comsol in bash to start Comsol 4.3b.

The first time I ran comsol server, it asked me to choose a username/password combo, so I made it the same as the combo used to log on to megatron.

Edit: I've also added a ~/.screenrc on megatron (based on this Stackoverflow answer) so that I don't constantly go nuts trying to figure out if I'm already inside a screen session.

  9129   Mon Sep 16 10:57:11 2013 SteveUpdateVACRGA scan at day 41

Quote:

 Valve configuration: Vacuum Normal

 

 Finally we reached 1.0e-5 Torr, cc1 (h) at day 41. There must be a leak or cold cathode gauge is not reading correctly. The gauge is new. It should not take this long.

All chamber annulos pressures are normal.

Attachment 1: pd76m41dRGA.png
pd76m41dRGA.png
Attachment 2: pd76m41d.png
pd76m41d.png
  9128   Fri Sep 13 19:22:01 2013 MasayukiUpdateLSCMICH calbration

 

 I made sure the yesterday's result was collect. I measured not only the error signal but also the feedback signal. And I compared those signals and measured the TF in order to confirm my servo filter model is not wrong.

The reason of dip at low frequency region is maybe the coherence of the ground motion. The ITMX and ITMY suspensions are put close. If ground motion has coherence, the mirrors move in common mode. That will suppress the free running noise. The attachment is the free running noise of Sep 13rd and Sep 12nd.

Attachment 1: noise.PNG
noise.PNG
  9127   Thu Sep 12 23:36:25 2013 MasayukiUpdateLSCMICH calbration

For Modelling of the OLTF, I measured the response of the BS suspension. I used the OSEM sensor for measurement. The attatchment1 is the measured TF from C1:SUS-BS_LSC_EXC to C1:SUS-BS_SUSPOS_IN1 with exciting with random force. The measured data was fitted and the resonant frequency is 1.029(±0.005) Hz and quality factor is 12.25 (± 0.2).  Additionally I did same measurement for ITMX and ITMY. The attachment 2 and 3 are the results for ITMX and ITMY. Each eigenfrequency and Q are 1.063 (±0.008) Hz and 7.33 (±0.13) (ITMX), 1.022 (±0.005) Hz and 9.41 (±0.09) (ITMY).

 After that, I locked the MICH with AS55, and measured the PSD of error signal. I compensated the that PSD by the modelled OLTF with this suspension TF and the servo TF. The result is in attachment 4. Above 1 Hz it is quite close to the previous data by Keiko (elog#6385) But below 1 Hz there is a large dip. The error signal has also this dip. I looked for a integral filter between 0.2 Hz and 1 Hz, but I connot find a such filter. And when I locked MICH with using ASDC, there was same dip at same frequency. I don't think it's true free running noise, and I will try to fix it.

I completely forgot to mention that I fitted the modelled OLTF into the measured OLTF. I used the fitted OLTF for compensation. 

 

 

Attachment 1: BSsus.PNG
BSsus.PNG
Attachment 2: ITMXsus.PNG
ITMXsus.PNG
Attachment 3: ITMY.PNG
ITMY.PNG
Attachment 4: free_running.PNG
free_running.PNG
  9126   Thu Sep 12 01:06:09 2013 JenneUpdateASCSRCL ASS implemented

I have modified the ASS model to also have an ASS for SRCL.  The input options are POPDC, POP110, AS110.  I suppose I could/should have included ASDC.

Screens are modified / made.  I haven't finished setting the servo gains and oscillator amplitudes, and all that jazz yet. 

Using the parameters that Koji had in elog 9116, I was able to get nice long DRMI locks (several on a ~10 minute time scale). 

I tried some pseudo-ANDing for the triggers, to no avail.  I was trying to have the trigger matrix row for the SRCL loop have 1*POP22 and 0.02*AS110, where the 0.02 is to scale AS110 so that it has a similar amplitude to POP22.  I then set threshold levels to ~250 for up, and 100 for down (I tried several different values for the up threshold).  I was watching the TRIG_MON_FAST channels for both PRCL and SRCL, and I wasn't able to get SRCL to be triggered only at the same times as PRCL using this technique.  Since we can get the DRMI to lock, perhaps my AND logic for the triggers is a low priority, but I think we'll need something like that if we want real logic.

 

  9125   Thu Sep 12 00:07:26 2013 MasayukiSummaryGreen LockingALS locking in both arms

Quote:

What was the beat freq for each arm?
The HF noise level depends on the frequency of the beat note.
As the BBPD has the freq dependent noise level. (See this entry)

 I'm not sure about the actual number of the beat frequency, but the beat frequency was almost same in both arms. And I took this measurement sometimes with slightly different beat frequency but the noise level didn't change so much.

  9124   Wed Sep 11 23:43:10 2013 KojiSummaryGreen LockingALS locking in both arms

What was the beat freq for each arm?
The HF noise level depends on the frequency of the beat note.
As the BBPD has the freq dependent noise level. (See this entry)

  9123   Wed Sep 11 23:34:37 2013 MasayukiSummaryGreen LockingALS locking in both arms

[manasa, masayuki]

We locked the XARM and YARM with using ALS control loop and we succeeded to lock stably both arms. The performance of the ALS was tested with a measurement of the calibrated error signal. (attachment 1)

- red and blue : the in-loop noise of ALS of each arm.

- green and purple:Stability of the beat-note frequency with the MC and the arm freely running.

Discussion

In the high frequency region, YARM has larger noise than XARM, and these noises were not there in previous measurements by Koji and Manasa (elog8865). You can see that in both of in-loop noise and free running noise. These noises may be caused by the Green PDH servo or hte phase tracker servo or any other electrical staff. We will start noise budget of these servo.

At higher frequency than UGF of ASL control loop, the loop does not suppress the noises at all, but the inloop and free running noise are not equivalent. I have no idea about that so far.

 

 

 

Attachment 1: noise.pdf
noise.pdf
  9122   Wed Sep 11 17:35:38 2013 JenneUpdateLSCALS requirement

I have done a quickie look at Optickle to see how the linewidth of an arm cavity changes versus the configuration. 

To do this, I make different configurations, and do a sweep of ETMX.  For each configuration, I find the max peak value, and then find the points that are at half that value.  The distance between them is the full width at half max.

I get:

FWHM_DRFPMI = 3.8750e-11  meters

FWHM_PRFPMI = 3.8000e-11  meters

FWHM_SRFPMI = 2.3200e-09  meters

FWHM_FPMI =   1.1900e-09  meters

So, for the ALS to hold within 1/10th of a linewidth for the full IFO configuration, we want the ALS noise to be on the order of 3 picometers RMS.  If I recall correctly, that's about an order of magnitude better than we currently have.

 

ArmLinewidthComparison.png

                 use LOG y-scale

EDIT 8 Nov 2013, JCD:  New log-y plot:

LinewidthComparison.png

  9121   Tue Sep 10 17:35:50 2013 Masayuki, ManasaUpdateLSCMICH calbration

Quote:

[Manasa, Masayuki]

We took a bunch of measurements. Transfer function and power spectrum using DTT. They will be used to obtain calibrated MICH in-loop and free-running noise. Detail Elog with plots will follow very soon.

 [Masayuki, Manasa]

Estimation of free-running MICH displacement noise:

Method 1. Assuming AS55_Q_err to be a linear sensor, as shown in (1) of figure below, free-running MICH noise (V_d) can be estimated by measuring V_err and the OLTF G. H can be estimated by using method explained in elog

 Method 2. Considering that the AS55_Q signal might be distorted or saturated, method 1 may not be precise. In method 2, we will use the ASDC as the sensor (S' in (3)) instead and lock MICH using ASDC in mid-fringe to calibrate the ITM actuators.

Figure:1

Schematic:

MICH_calib_loops.png

What we did:

1. Estimate H' from free-running ASDC signal (bright to dark fringe).
2. With MICH locked on ASDC, give an excitation signal to C1:LSC-SUS_XXXX_EXC (XXXX could be ITMX or ITMY) and measure R'. [(3) of schematic]
3. Measure OLTF of MICH locked on ASDC (hence estimate L). [(3) of schematic]
4. With MICH locked on AS55_Q, give an excitation signal to C1:LSC-SUS_XXXX_EXC (XXXX could be ITMX or ITMY) and measure R1. [(2) of the schematic] 

Results/Plots: 

Figure:2

OLTF of MICH locked on ASDC

 OLTF_MICHDC.png

 

Figure2:

Actuator excitation to MICH transfer function (MICH locked using ASDC) 

MICH_DC_resp.png

* y axis (no units)

Figure 3:
Actuator excitation to MICH transfer function (MICH locked using AS55Q)

MICH_RF_resp.png 

* y axis (no units)

Figure 4:
Free-running MICH noise

MICH_free_noise.png 

Discussion: 

1. By using the second sensor, we also eliminate the effect of the MICH servo loop locked on AS55_Q (Estimated V_d does not depend on G but only on G').

2. The free-running MICH noise is still suppressed at 1Hz. This should be coming from the effect of the UGF of the loop at ~10Hz and the vicinity to the pendulum frequency at 1Hz.

 

Edit/Masayuki// This noise curve is not collect, especially in low frequency region. We used the measured OLTF for compensating the free running noise, but that is not collect in low frequency region. So we should model the OLTF and fit that into the measured OLTF. We will fix this soon.

 

 

  9120   Tue Sep 10 15:43:01 2013 SteveUpdateVACTP3' dry pump is replaced

Quote:

 

 TP3 foreline's dry pump is getting noisier and noisier.  Turbo TP3 is pumping on the annulos. The foreline pressure is 7.2 mTorr and it is not degrading. It was swapped in March 5, 2013

The seal is very good, but the bearing is dying.

 

 

 

 The drypump is replaced at 95,781 hrs on TP3 controller time. The foreline pressure is 30 mTorr and dropping.

 It is 13 mTorr after 17 hours of pumping.

  9119   Tue Sep 10 09:18:22 2013 SteveUpdateVACdry pump is getting loud

 

 TP3 foreline's dry pump is getting noisier and noisier.  Turbo TP3 is pumping on the annulos. The foreline pressure is 7.2 mTorr and it is not degrading. It was swapped in March 5, 2013

The seal is very good, but the bearing is dying.

 

 

 

  9118   Mon Sep 9 20:46:28 2013 MasayukiUpdateLSCMICH calbration

[Manasa, Masayuki]

We took a bunch of measurements. Transfer function and power spectrum using DTT. They will be used to obtain calibrated MICH in-loop and free-running noise. Detail Elog with plots will follow very soon.

  9117   Mon Sep 9 15:33:06 2013 SteveUpdateVACvertex crane repair is scheduled

Quote:

 

 [Fred Goldbar, Mike Gerfen, Dennis Coyne and Steve]

We inspected the hinge, 1.25" cross pin and I-beams. It is hard to explain what is  causing the folding I-beam corner to jam against the main I-beam.

To limit the motion of the folding I-beam  cross pin bushing will be added. This will take a week to complete.

 

 

 

 

 

 KoneCrane contact John McDaniel (562) 903 - 1371,

Wednesday, September 18,  folding I-beam will be removed.  KoneCrane will start working at 7:30am and they should be out by 12:30pm

 Friday,  September 20, reinstalling machined hinge on the I-beam. Same timing schedule as Wednesday.

  9116   Fri Sep 6 23:01:08 2013 KojiUpdateLSCStable DRMI lock was recovered from the impact on the RF system modification

Summary

Stable DRMI lock was recovered. The AS110 phase was adjusted. PRCL and MICH were locked with REFL33I and REFL165Q.
Still SRCL is controlled with REFL55Q.


PRMI sensing matrix

Thursday night, Jenne and I found DRMI can not be locked at all. Also the PRMI lock with REFL55 showed change in the optical gain.

In order to investigate what is happening, the PRMI sensing matrix was measured and compared with the previous one taken in the night of 8/26.

SensMat_PRMI_1000cts_580Hz_2013-08-26_235635.pngVSSensMat_PRMI_1000cts_580Hz_2013-09-06_201137.png

It shows that some signals are unchanged, some are partial change, and some are completely different.
My intuition saids something is wierd with the sensing matrix measurement.
Right now I can't trust these plots.

- Jenne and I have adjusted REFL55 demod angle so that REFL55Q has no PRCL. And I have confirmed with DTT that this is still true.
  However, the radar chart shows that REFL55Q is almost correct phase for PRCL instead of MICH.

- REFL11 shows the same amplitude and angle as before. But POX11/POY11 shows different MICH angle.

- I have rotated REFL55 demod phase and remearsured the sensing matrix. Evrything else looked same but REFL55.
  Since REFL55I&Q were not used for the control for this measurement, what we expect is to see no change of the sensing matrix and
  only see the angle of "I"&"Q" rotates. But the result was different from the expectation.

DRMI locking

Since no real info was obtained from the sensing matrix, I had to make a fight without any weapon.
After sevral hours of work, stable DRMI lock was recovered.

Basically I gave larger gains to REFL55 signals: REFL55I for SRCL was 100 instead of 1, and REFL55Q for MICH was 2 instead of 0.1.
This was enough to get a second locking. Using this short sections, I have optimized the FM triggers and the gain boosts (i.e. FM1)
as well as the mirror alignment.

Then, PRM ASS was left running during the lock. This actually stabilized the lock a lot.
This made thee lock indefinite.

The demod phase of AS110I was adjusted so that AS110Q fluctuates around zero.
In this condition, the nominal AS110I was 7300 with the whitening gain of 30dB.

Note that the AS110I&Q were also measured with PRMI. With the same phase and gains, AS110I and Q were -35,  -170, respectively.
Do we expect to have this phase shift? If I believe these numbers, the aplitude of 110MHz at the optimal phase is 173,
The ratio of AS110 between DRMI and PRMI is 7300/173 = 42. This corresponds to the ratio of the 110MHz sideband power at the AS port.
According to the wiki, this ratio shoud be ~160.

AS110I was in fact glitchy as you can see in the StripTool chart. I wonder this signal is suitable for the normalization or not.


=== SENSING ===

REFL11 -67deg / whitening gain 0dB
REFL33 -20deg / whitening gain 30dB
REFL55 45deg / whitening gain 6dB
REFL165 96deg / whitening gain 45dB

POP110 69deg whitening on / 15dB
POP22 102.2deg whitening on / 21dB
AS110 145deg whitening off / 30dB (seems to be related to AS11 whitening setting)

=== INPUT MATRIX ===

REFL11I x -0.125 => PRCL (REFL33I x 2.5 was also OK)
REFL55I x 100 => SRCL
REFL55Q x 2 => MICH (REFL165Q x 0.1 was also OK)

=== NORMALIZATION / TRIGGER ===

No normalization

Trigger settings
MICH POP22I UP:50 DOWN:10
PRCL POP22I UP:50 DOWN:10
SRCL POP22I UP:50 DOWN:25

=== SERVO FILTERS ===

MICH x -0.8 FM4/5 ON, no limitter
FM Trigger: delay 2sec, FM1 (modified from 6dB to 20dB), FM2, FM3

PRCL x +0.035 FM4/5 ON, no limitter
FM Trigger: delay 0.5sec, FM2/3/6

SRCL x -0.1 FM4/5 ON, no limitter
FM Trigger: delay 5sec, FM1, FM2

=== OUTPUT FILTERS ===

MICH => PRM -0.267 / BS +0.5

PRCL => PRM +1.0

SRCL => SRM +1.0

=== VIOLIN FILTER TRIGGER ===

delay 1sec: FM1/FM2/FM3/FM6

=== ASC/ASS ===

PRM ASC UP:50 DOWN:25
PITCH&YAW: FM1/9 (ALWAYS ON) + FM2/3 (turned on by the up-script)

PRM ASS left turned on for slow tracking

Attachment 1: DRMI.png
DRMI.png
  9115   Fri Sep 6 09:27:10 2013 SteveUpdateVAC31 days after pumpdown

Quote:

 Valve configuration: Vacuum Normal

 

 

Attachment 1: day31vac.png
day31vac.png
  9114   Thu Sep 5 21:07:09 2013 JenneUpdateLSCStarted work on logic for triggering

I want something like an "AND" for the degree of freedom triggers.  Koji and I talked through an idea, and I have it running in the c1tst model, but the logic isn't working like I expect, so I need to look into it more before I can put it into the lsc model.

ELOG V3.1.3-